OpenStudy (anonymous):

Solve by factoring. 4x2 + 23x + 15 = 0

OpenStudy (anonymous):

ooh i like these

OpenStudy (freckles):

\[ax^2+bx+c=0\] So what you want to do is find factors of a*c that have product a*c and sum b.

OpenStudy (anonymous):

(x+5)*(4*x+3)

OpenStudy (anonymous):

Good! Can you help me?

OpenStudy (freckles):

find two factors*

OpenStudy (freckles):

So what is a here? And what is b here? And what is c here?

OpenStudy (anonymous):

\[4x ^{2}+23x+15=0\]

OpenStudy (anonymous):

(x+5)*(4*x+3) x= -5 and -3/4 right

OpenStudy (freckles):

@timo86m I don't think it helps to give the answer. Can you explain how you got those factors?

OpenStudy (anonymous):

A:x = 0.8 or x = -0.75 C:x = 0.75 or x = 5 B:x = -3 or x = -1.25 D:x = -0.75 or x = -5

OpenStudy (freckles):

@Makayla43 if you want to learn this, I'm trying to teach you.

OpenStudy (freckles):

What is a,b, and c in your equation?

OpenStudy (anonymous):

@KNicksfan4life :D (:

OpenStudy (anonymous):

(4x-3)(x-5)

OpenStudy (anonymous):

@freckles what do you mean a,b,and c?

OpenStudy (anonymous):

is that rite

OpenStudy (anonymous):

it not right

OpenStudy (freckles):

\[ax^2+bx+c=0\] \[4x^2+23x+15=0\]

OpenStudy (anonymous):

(x+5)*(4*x+3) if you not sure you can always expand this

OpenStudy (anonymous):

Solve by factoring. -3x2 - x + 2 = 0

OpenStudy (anonymous):

-(x+1)*(3*x-2)

OpenStudy (anonymous):

\[-3x ^{2}-x+2=0\]

OpenStudy (anonymous):

-3x2 - x + 2 = 0 you actually mean -3x^2 - x + 2 = 0

OpenStudy (anonymous):

A:x = 0.5 or x = 0.33333 C:x = 1 or x = -0.66667 B:x = 0.33333 or x = -2 D:x = -1 or x = 0.66667

OpenStudy (anonymous):

I think it d

myininaya (myininaya):

You guys how is this teaching and how is this learning? Please try to teach @Makayla43 and not just give the answer. And @Makayla43 you shouldn't be just looking for answers. I think freckles here was trying to teach you how to do this.

OpenStudy (anonymous):

Omg.

OpenStudy (anonymous):

i teached, I said expand my factor if she doubts it lol.

myininaya (myininaya):

But @timo86m how do you get the factors? Please explain that part so she will know how to do it.

OpenStudy (anonymous):

Goodbye. He sent me a message a explained.

OpenStudy (anonymous):

4x2 + 23x + 15 = 0 number that up to 15 and product is 23

OpenStudy (anonymous):

To factor is mostly guess work that is where I draw the line :)

OpenStudy (anonymous):

number that up to c and product is b coefficient. It is more important to show why it works then how to do something that is mostly guess work.